A flea jumps straight up to a maximum height of 0.500 m

m. What is its initial velocity v


v0v_0 as it leaves the ground?


1
Expert's answer
2021-09-26T08:39:45-0400

Given:

"h_{\max}=0.500\:\rm m"


The maximum height

"h_{\max}=\frac{v_0^2}{2g}"

Hence, the initial velocity of a flea

"v_0=\sqrt{2gh_{\max}}=\sqrt{2*9.81*0.500}=3.13\:\rm m/s"